A. If you want to earn 9% annual simple interest on an investment, how much should you pay for a note that will be worth $24,400.00 in 2 months.

B.If $9,000.00 is invested at 16% annual simple interest, how long does it take to be worth $37,800.00.

It will take ____years? (round to the nearest whole).
Answer #1

1.

Using the simple interest formula

I = PRT----------(i)

where

I +P = 24400

I = 24400-P

r = 9% =0.09

t = 2/12 = 0.167


putting values in (i)

24400 - P = P(0.09)(0.167)
24400 - P = 0.01503P
24400 = 1.01503P
P = 24400/1.01503 = 24038.69

Amount to be paid is $24,038

2.

I = PRT

37800-9000 = 9000x0.16xT

T = 28800/1440

T = 20 years